# Entropy-Enhanced Fractional Quantum Anomalous Hall Effect

### v1

@@ Abstract
Introduction.— The fractional quantum anomalous Hall effect [1, 2] (FQAH) is the zero magnetic field analog of the celebrated fractional quantum Hall effect [3, 4]. Recent advances in moiré materials have recently enabled experimental observation of this topological strongly-correlated phase of matter in twisted transition metal dichalcogenides [5, 6, 7, 8], and crystalline graphene/hBn moiré superlattices [9, 10, 11].
@@
Figure 1: Schematic description of the phase diagram emerging from our model. The ground state of the system is tuned by some parameter, e.g., displacement field $D$, determining the energy difference between the fractional quantum anomalous Hall phase (FQAH) on the left side, and the extended integer quantum anomalous Hall phase (EIQAH) on the right. At higher temperatures, the phase boundary (dark blue solid line) shifts as $T\propto\sqrt{\delta u}$ [see Eq. (4)], favoring the phase with higher edge-state entropy. As the linear dimension of the system $L$ increases, the shift of the boundary becomes more subtle (red dotted line) and the entropy driven transition may be preempted by a transition to a metallic phase.
@@
We provide key predictions which may confirm or invalidate the proposed scenario. Namely, under our assumptions the entropic effect is necessarily device-size-dependent, and should be virtually undetectable for a large enough sample. Moreover, the temperature dependence of the phase boundary between the phases should follow $T\propto\sqrt{\delta u}$, with $\delta u$ the ground state energy difference between the phases. Finally, observation of the EIQAH-to-FQAH crossover at various values of electric displacement fields suggests that this $\delta u$ should be relatively insensitive to displacement field changes.
Theory.— We aim to provide a phenomenological characterization of the competition between the extended quantum anomalous Hall phase ($I$) and the fractional quantum anomalous Hall phase ($F$), via a description of the free energy associated with these phases. Assuming a temperature scale far below the respective transition temperatures of these phases to a metallic phase, we approximate the free energies as,
@@
Let us denote by $\Omega$ and $\mathtt{P}$ the area and perimeter of the device. The entropy density of the phase $\mu$ originating in its linearly dispersing edge states is
@@
where $v_{i}$ are the edge mode velocities, and $k_{B}$ is the Boltzmann constant. It is then useful to define the potential energy density difference $\delta u\equiv\left(U_{F}-U_{I}\right)/\Omega$, and the velocity scale
@@
The transition between the two phases, described by Eq. (1), is then given by the condition,
@@
Taking into account realistic parameters from the experiment in Ref. [12], one may arrive at a rough estimate of $\delta u$ given Eq. (4) characterizes the transition. Considering an electronic density of $n_{e}\approx 0.5$ 1012 cm-2, the difference in energies between the fractional and extended integer phase $\delta u/\left(k_{B}n_{e}\right)$ is on the order of $O$(5 mk / electron).
@@
This point is demonstrated in Fig. 2, which examines a key quantum geometrical property of the lowest-lying conduction band in pentalyer graphene, namely the trace of the Fubini-Study metric ${\rm tr}g$. This quantity is believed to play a key role in stabilizing the fractional quantum anomalous Hall phase in a given band [27, 28, 29, 30, 31, 32, 28, 30, 33, 34, 35, 36]. For simplicity, we omit the effect of the moiré-inducing hBN layer, since electrons are mostly polarized to the graphene layer far away from it in the relevant regime. A change of 5% to the potential difference between the outermost graphene layers hardly has any effect, as seen by comparing Fig. 2a–b. In Fig. 2c we plot the relative change in this quantity, which remains much smaller than the relative change in $D$ throughout the majority of the enclosed Fermi volume at the relevant densities.
@@
If the excess edge density of states ($\propto 1/v^{*}$) is large enough, or the potential energy difference $\delta u$ is rather small, the EIQAH-to-FQAH “melting” can be observed for realistic device sizes.
We note that in the presence of disorder, specifically of the kind that couples differently to the $F$ and $I$ phases, the impact of the edge-state entropy may be further enhanced. Domain walls separating the topologically distinct phases contain protected edge modes which carry entropy at finite temperature. A transition from a uniform $I$ phase at low-temperatures to a domain-patterned phase at higher temperature, analogous to the so-called Chern mosaic [37], is expected in such a scenario. The interplay between disorder and domain wall entropy driven transitions has been thoroughly explored in Ref. [38]. The signature this sort of transition would have on transport is not universal, would depend on the position of the domains relative to the contacts, and would thus be hard to reproduce.
Discussion.— We have presented a scenario accounting for the emergence of FQAH upon heating of an EIQAH ground state. The culprit in this scenario is the excess entropy of the FQAH edge states. If the ground-state energy difference between the phases is sufficiently small, the system may gain energy by “partially melting” in its edges. This is enabled by transitioning into the intermediate FQAH, which allows additional gapless excitations on its edge.
@@
Acknowledgements.
GS acknowledges enlightening discussions with Gil Refael, as well as support from the Walter Burke Institute for Theoretical Physics at Caltech, and from the Yad Hanadiv Foundation through the Rothschild fellowship.

### v2

@@ Abstract
Introduction. The fractional quantum anomalous Hall effect [1, 2] (FQAH) is the zero magnetic field analog of the celebrated fractional quantum Hall effect [3, 4]. Recent advances in moiré materials have recently enabled experimental observation of this topological strongly-correlated phase of matter in twisted transition metal dichalcogenides [5, 6, 7, 8], and crystalline graphene/hBn moiré superlattices [9, 10, 11].
@@
Figure 1: Schematic description of the phase diagram emerging from our model. The ground state of the system is tuned by some parameter, e.g., displacement field $D$, determining the energy difference between the fractional quantum anomalous Hall phase (FQAH) on the left side, and the extended integer quantum anomalous Hall phase (EIQAH) on the right. The left and right insets illustrate a possible edge picture of the respective phases. The integer EIQAH phase a single chiral edge mode, whereas the fractional phase may host a rich plethora of edge modes, each contributing to the overall entropy. At higher temperatures, the phase boundary (dark blue solid line) shifts as $T\propto\sqrt{\delta u}$ [see Eq. (4)], favoring the phase with higher edge-state entropy. As the linear dimension of the system $L$ increases, the shift of the boundary becomes more subtle (red dotted line) and the entropy driven transition may be preempted by a transition to a metallic phase.
@@
We provide key predictions which may confirm or invalidate the proposed scenario. Namely, under our assumptions the entropic effect is necessarily device-size-dependent, and should be virtually undetectable for a large enough sample. Moreover, the temperature dependence of the phase boundary between the phases should follow $T\propto\sqrt{\delta u}$, with $\delta u$ the ground state energy difference between the phases. Finally, the appreciable evolution of the EIQAH-FQAH boundary as a function of electric displacement field and temperature suggests that this $\delta u$ should be relatively insensitive to displacement field changes.
Theory. We aim to provide a phenomenological characterization of the competition between the extended quantum anomalous Hall phase ($I$) and the fractional quantum anomalous Hall phase ($F$), via a description of the free energy associated with these phases. Assuming a temperature scale far below the respective transition temperatures of these phases to a metallic phase, we approximate the free energies as,
@@
Let us denote by $\Omega$ and $\mathtt{P}$ the area and perimeter of the device. We will define the potential energy density difference between the two phases $\delta u\equiv\left(U_{F}-U_{I}\right)/\Omega$, which is a phenomenological parameter of our theory. It is determined by details of the underlying microscopic theory of both the fractional and integer phases, which are beyond the scope of this work. The entropy density of the phase $\mu$ originating in its linearly dispersing edge states is
@@
where $v_{i}$ are the edge mode velocities in the phase $\mu$, and $k_{B}$ is the Boltzmann constant. It is then useful to define the velocity scale
@@
The transition between the two phases, which is of first-order by construction (we consider the competition between the $G_{\mu}$ described in Eq. (1), is then given by the condition [26],
@@
Taking into account realistic parameters from the experiment in Ref. [12], one may arrive at a rough estimate of $\delta u$ given Eq. (4) characterizes the transition. Considering an electronic density of $n_{e}\approx 0.5$ 1012 cm-2, and parameters $v^{*}\sim 10^{3}$ m/sec, $\tilde{T}\sim 0.1$ K, $L\sim 1$ $\mu$m, the difference in energies between the fractional and extended integer phase $\delta u/\left(k_{B}n_{e}\right)$ is on the order of $O$(5 mk / electron).
@@
An in-depth study of the effect of $D$ on the stability of the various FQAH states, and on its energetic competition with the EIQAH phase (whose nature is yet unresolved) is beyond the scope of this work. Such a study requires of course, among other things, consideration of the electron-electron interactions necessary to facilitating a fractional phase. Let us, however, examine a single-particle property which has been shown to play a key role in stabilizing the fractional quantum anomalous Hall phase in a given band, namely the trace of the Fubini-Study metric ${\rm tr}g$ [27, 28, 29, 30, 31, 32, 28, 30, 33, 34, 35, 36]. For simplicity, we omit the effect of the moiré-inducing hBN layer, since electrons are mostly polarized to the graphene layer far away from it in the relevant regime. A change of 5% to the potential difference between the outermost graphene layers hardly has any effect on this quantum geometric property, as seen by comparing Fig. 2a–b. In Fig. 2c we plot the relative change of this metric, which remains much smaller than the relative change in $D$ throughout the majority of the enclosed Fermi volume at the relevant densities. This simple calculation demonstrates that at the experimentally relevant values of displacement field, a decisive parameter controlling the FQAH physics remains quite insensitive to changes in $D$. Such insensitivity is this consistent with a very small $\delta\mu_{D}$.
@@
Using the estimate for $\delta u$ we obtained above, given $T_{c}$ which is of order of a few hundred mK, one may estimate $\tilde{L}\sim O\left(10\,\mu{\rm m}\right)$. If the excess edge density of states ($\propto 1/v^{*}$) is large enough, or the potential energy difference $\delta u$ is rather small, the EIQAH-to-FQAH “melting” can be observed for realistic device sizes.
We note that in the presence of disorder, specifically of the kind that couples differently to the $F$ and $I$ phases, the impact of the edge-state entropy may be further enhanced. Domain walls separating the topologically distinct phases contain protected edge modes which carry entropy at finite temperature. A transition from a uniform $I$ phase at low-temperatures to a domain-patterned phase at higher temperature, analogous to the so-called Chern mosaic [37], is expected in such a scenario. In such a mosaic phase, $\alpha_{\rm geo.}$ effectively becomes much smaller, since one replaces the perimeter of the device $\tt P$, with the overall length of domain walls in the system, roughly proportional to $\tt P$ times the number of domains. The interplay between disorder and domain wall entropy driven transitions has been thoroughly explored in Ref. [38]. The signature this sort of transition would have on transport is not universal, would depend on the position of the domains relative to the contacts, and would thus be hard to reproduce.
Discussion. We have presented a scenario accounting for the emergence of FQAH upon heating of an EIQAH ground state. The culprit in this scenario is the excess entropy of the FQAH edge states. If the ground-state energy difference between the phases is sufficiently small, the system may gain energy by “partially melting” in its edges. This is enabled by transitioning into the intermediate FQAH, which allows additional gapless excitations on its edge.
@@
G.S. acknowledges enlightening discussions with Gil Refael, as well as support from the Walter Burke Institute for Theoretical Physics at Caltech, and from the Yad Hanadiv Foundation through the Rothschild fellowship.